OpenStudy (anonymous):

Can someone help please. A gas can holds 2.9 gal of gasoline.What is this quantity in cubic centimeters?

OpenStudy (jhannybean):

\[\sf 2.9~gal~\times~\frac{3.785~\times~10^3~cm^3}{1~gal}=~?~~cm^3\]Your answer will be 2 sig figs, so use scientific notation.

OpenStudy (anonymous):